OpenStudy (anonymous):

Which of the following has a slope of 2 and a y-intercept of 4? y = 2x + 4 y = 2x + 4 y = 2x 4 y = 2x 4

OpenStudy (lukecrayonz):

im assuming the ones that have no sign are subtraction, it just doesn't show :) but you have two of the same!

OpenStudy (lukecrayonz):

But the answer is y=2x+4, as y=mx+b, m = slope, b = y intercept :)

OpenStudy (anonymous):

lol sorry heres a better one Which of the following has a slope of 2 and a y-intercept of -4? (Points : 4) y = 2x + 4 y = 2x + 4 y = 2x -4 y = -2x - 4

OpenStudy (lukecrayonz):

Well which one do you think is the answer?

OpenStudy (anonymous):

Which of the following has a slope of 2 and a y-intercept of -4? (Points : 4) y = -2x + 4 y = 2x + 4 y = 2x -4 y = -2x - 4

OpenStudy (lukecrayonz):

What do you think the answer is? I don't want to just give you answers!

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

but not sure

OpenStudy (anonymous):

t hat why im asking

OpenStudy (lukecrayonz):

Well no, look at the formula y=mx+b, the formula for a line. in that formula, m is your slope, b is your y intercept. So a slope of 2, y=2x so far, and your y intercept of -4, y=2x-4! :)